OpenStudy (anonymous):

At Ron's Roller Rink, the number of customers has been decreasing at a steady rate of 5% per year. If there were 900 skaters per week in 2000, what is a good estimate for the number of skaters per week in 2006? A. 778 skaters B. 662 skaters C. 695 skaters D. 732 skaters

OpenStudy (anonymous):

900 in 2000, x in 2006 6 periods, -5% per period. Each period has 95% of the skaters the last period had. number in 2001 = number in 2000 * 0.95 number in 2002 = number in 2001 * 0.95 Substituting, number in 2002 = number in 2000 * 0.95 * 0.95 If you continue this, you will notice a pattern. After n periods, the amount will be initial * percent each time to the power of n. Pluggin values in: skaters in 2006 = 900 * 0.95^6 skaters in 2006 approx = 661.5827 Answer is B. 662.

OpenStudy (anonymous):

Lol, defined x and never even used it. Disregard it, thought about explaining it a diff way but decided against it.

OpenStudy (anonymous):

Thanks